# Spinning masters: on the impact of tidal forces and protohalo size on early spin evolution

### v1

@@ 6.1 Amplitude
In Figure 9 we show the median evolution of the specific AM amplitude, both measured in the simulation and predicted by our implementation of the TTT. The gray solid lines represent the median of the AM distribution of protohalos at each snapshot, whereas the surrounding gray shaded regions indicate the corresponding interquartile range. The black dotted lines show the expected behavior according to Eq. (4), with a smoothing scale $R_{\mathrm{s}}=R_{\mathrm{eq}}$ for each protohalo. From left to right, each panel corresponds to a subsample of protohalos of increasing mass, as indicated by the top-left labels. The shaded region in the righ-hand side of each panel indicates the final, highly non-linear stage, where the TTT is not expected to be valid.
@@
As expected, particles residing in the inner regions of the halo at present time exhibit lower AM at late times compared to those in the outer regions. This trend becomes particularly evident during the final stages of evolution, where the specific AM of these particles decreases sharply. This sharp decline indicates a net loss of AM as the particles settle into the inner regions of the protohalos.
Interestingly, the earlier stages of evolution reveal a more complex picture. Despite their current location in the inner halo, these particles initially had comparable or even higher specific AM than those located in the outer regions of the protohalo. This result suggests that particles contributing to the protohalo span a broad range of specific AM values and that significant redistribution of AM occurs throughout the halo’s assembly process. By the end of the simulation, particles that lose a substantial fraction of their specific AM to the outer regions tend to migrate inward, settling into the halo’s core.
@@ 6.3 Mass dependent Spin-Flip
An additional key result is that the AM direction of the innermost particles at $z=0$ does not display the well-known mass-dependent “spin flip” trend observed in full halos (Aragón-Calvo et al., 2007b; Hahn et al., 2007; Codis et al., 2015; Ganeshaiah Veena et al., 2018, 2021b). For the total halo spin, it is typically found that low-mass halos show an alignment with $\hat{\mathbf{e}}_{3}$, the filament spine, while high-mass halos tend to exhibit perpendicular AM alignments. In contrast, for particles at $0.1R_{\mathrm{vir}}$, we find that the AM is perpendicular to $\hat{\mathbf{e}}_{3}$ regardless of the halo mass.
@@ 7 Summary and conclusions
We have evaluated the ability of the TTT to predict both the amplitude and direction of protohalo AM at high redshift. Our results show that the accuracy of TTT predictions depends sensitively on the smoothing scale used to compute the tidal field. At fixed smoothing lengths, smaller scales yield more accurate predictions for both the amplitude and orientation of the AM, particularly for low-mass protohalos. However, the optimal filter scale increases with protohalo mass, reflecting the coupling of larger protohalos to longer-wavelength tidal modes. Crucially, when the filter scale is adjusted to the equivalent radius $R_{\mathrm{eq}}$ of each protohalo, the model achieves its best overall performance.
@@
In terms of direction, our results demonstrate a systematic and mass-dependent evolution of the AM orientation relative to the cosmic web. Early alignments with the spine of present-day filaments ($\hat{\mathbf{e}}_{3}$) degrade over time, particularly due to a gradual rotation of the AM vector around the intermediate filament axis ($\hat{\mathbf{e}}_{2}$), effectively transferring alignment toward the principal axis of collapse ($\hat{\mathbf{e}}_{1}$). This transition is most evident during the quasi-linear regime ($z\sim 10$ to $z\sim 1$), after which chaotic evolution in the non-linear regime tends to randomize AM directions, although some coherence persists, especially for massive halos. Notably, the final alignments at $z=0$ appear more strongly shaped by the quasi-linear evolution than by the chaotic transformations of the non-linear stage. These findings suggest that the breakdown of TTT is not abrupt but unfolds progressively, driven by the combined effect of internal dynamical processes, the differential coupling with the large-scale tidal field and the evolving structure of the cosmic web.
@@
Altogether, our results show that both the depth of particles within halos and the time evolution of the protohalo shape critically influence the AM direction. Early quasi-linear torques can generate strong alignments even for initially weakly aligned regions, but these are progressively erased by nonlinear processes. Properly modeling the inertia tensor evolution is thus essential to extend TTT predictions beyond the initial linear regime.
@@ Acknowledgements
This work has been partially supported by the Argentine Consejo Nacional de Investigaciones Científicas y Técnicas (CONICET). This work used computational resources from UNC Supercómputo (CCAD) – Universidad Nacional de Córdoba (https://supercomputo.unc.edu.ar), which are part of SNCAD, República Argentina.
@@ Appendix A Cosmic web identification with NEXUS+
The NEXUS+ suite of cosmic web identifiers represents an elaboration and extension of the original Multiscale Morphology Filter (Aragón-Calvo et al., 2007b, a, 2010) algorithm and was developed with the goal of obtaining a more physically motivated and robust method (for short summaries see also Libeskind et al. (2018); Kugel & van de Weygaert (2024)). NEXUS+ is the principal representative of the full NEXUS+ suite of cosmic web identifiers (see Cautun et al., 2013). These include the options for multiscale morphology identifiers on the basis of the linear density, the logarithmic density, the velocity divergence, the velocity shear and tidal force field. NEXUS+ has incorporated these options in a versatile code for the analysis of cosmic web structure and dynamics following the realisation that they are significant physical influences in shaping the cosmic mass distribution into the complexity of the cosmic web.
@@ A.1 Hessian Geometry and Morphological Identity
The basic setup of MMF/NEXUS+ is that of defining a four-dimensional scale-space representation of the input field $f(\vec{x})$. In nearly all implementations this achieved by means of a Gaussian filtering of $f(\vec{x})$ over a set of scales $[R_{0},R_{1},...,R_{N}]$,
@@
[…]
Note that the definition for the Hessian includes the normalisation term $R_{n}^{2}$. The key element of the MMF/NEXUS+ formalism is the morphological information contained in the eigenvalues of the Hessian matrix, $h_{1}\leq h_{2}\leq h_{3}$. By applying a set of morphology filters on these scaled eigenvalues (see Aragón-Calvo et al., 2007a; Cautun et al., 2013) this is translated into a scale dependent environment signature $\mathcal{S}_{R_{n}}(\mathbf{x})$ that represents the geometry at the corresponding scale.
@@ A.2 Scale Space and Multiscale Structure
To analyse the multiscale nature of the cosmic web, the Scale-Space representation of the cosmic mass distributions produces a sequence of copies of the data having different resolutions (Florack et al., 1992; Lindeberg, 1994). At each location $\vec{x}$ in the probed volume, it involves an extra dimension, scale, that yields the eigenvalues of the Hessian filtered at the corresponding (Gaussian) scale and the scale dependent environment signature $\mathcal{S}_{R_{n}}(\mathbf{x})$.
A feature searching algorithm is applied to the combined set of scaled copies in order to identify the scale at which, locally, we find the strongest morphological signature. It involves the combination of the complete set of scale-dependent environmental signatures to find the maximum signature for all scales
@@ A.3 Signature & Versions
The final outcome of the MMF/NEXUS+ procedure is a field which at each location $\vec{x}$ specifies what the local morphological signature is, cluster node, filaments, wall or void. The resulting field $\delta^{\text{NEXUS}}_{j}(\mathbf{x})$ is zero when the volume is not identified as cosmic web element $j$ and is one when the volume elements is identified as element $j$. Here $j$ is either filaments, nodes or walls. In this identification we also intrinsically include the identification for voids which is defined as the volume elements that are neither a filament, node or wall.
@@ A.4 NEXUS+
NEXUS+ is the principal representative of the full NEXUS+ suite of cosmic web identifiers (see Cautun et al., 2013). It is the density field NEXUS+ version with the highest dynamic range. As input it takes a regularly sampled density field. In a first step, the input field is Gaussian smoothed using a $\log$Filter filter that is applied over a set of scales $[R_{0},R_{1},...,R_{N}]$, with $R_{n}=2^{n/2}R_{0}$. It produces the logarithmic density field
@@
The NEXUS+ algorithm performs the environment detection by applying the above steps first to knots, then to filaments and finally to walls. Each volume element is assigned a single environment characteristic by requiring that filament regions cannot be knots and that walls regions cannot be either knots or filaments. The remaining regions are classified as voids.

### v2

@@ 6.1 Amplitude
In Figure 9 we show the median evolution of the specific AM amplitude, both measured in the simulation and predicted by our implementation of the TTT. The gray solid lines represent the median of the AM distribution of protohalos at each snapshot, whereas the surrounding gray shaded regions indicate the corresponding interquartile range. The thick black dotted lines show the expected behavior according to Eq. (4), with a smoothing scale $R_{\mathrm{s}}=R_{\mathrm{eq}}$ for each protohalo. From left to right, each panel corresponds to a subsample of protohalos of increasing mass, as indicated by the top-left labels. The shaded region in the righ-hand side of each panel indicates the final, highly non-linear stage, where the TTT is not expected to be valid.
@@
As expected, particles residing in the inner regions of the halo at present time exhibit lower AM at late times compared to those in the outer regions. This trend becomes particularly evident during the final stages of evolution, where the specific AM of these particles decreases sharply. As illustrated by the thin dotted black curve, which represents the case of constant AM in these diagrams, this sharp decline indicates a net loss of AM as the particles settle into the inner regions of the protohalos.
Interestingly, the earlier stages of evolution reveal a more complex picture. Despite their current location in the inner halo, these particles initially had comparable or even higher specific AM than those located in the outer regions of the protohalo. In effect, Figure 9 shows that, from $z=80$ down to $z\sim 2$, the median specific AM of the particles that end up at $0.5R_{\mathrm{vir}}$ at $z=0$ (and, to a lesser extent, also those at $0.1R_{\mathrm{vir}}$) is higher than that of the total FOF halo, once all values are normalized by $a^{2}\dot{D}$. In the regime where both curves are roughly constant, one clearly lies above the other. For example, for low-mass halos, the median $\log(j/a^{2}\dot{D})$ for the $0.5R_{\mathrm{vir}}$ particles is $\sim-1.75$, whereas for the total FOF halo it is $\sim-1.95$. In linear terms, this implies that the typical specific AM of the inner particles is about $60\%$ higher than that of the full FOF protohalo.
This result suggests that particles contributing to the protohalo span a broad range of specific AM values and that significant redistribution of AM occurs throughout the halo’s assembly process. By the end of the simulation, particles that lose a substantial fraction of their specific AM to the outer regions tend to migrate inward, settling into the halo’s core.
@@ 6.3 Mass dependent Spin-Flip
An additional key result is that, at present time, the AM direction of the innermost particles does not display the well-known mass-dependent “spin flip” trend observed in full halos (Aragón-Calvo et al., 2007b; Hahn et al., 2007; Codis et al., 2015; Ganeshaiah Veena et al., 2018, 2021b). For the total halo spin, it is typically found that low-mass halos show an alignment with $\hat{\mathbf{e}}_{3}$, the filament spine, while high-mass halos tend to exhibit perpendicular AM alignments. In contrast, for particles at $0.1R_{\mathrm{vir}}$, we find that the AM is perpendicular to $\hat{\mathbf{e}}_{3}$ regardless of the halo mass.
@@ 7 Summary and conclusions
We have evaluated the ability of the TTT to predict both the amplitude and direction of protohalo AM at high redshift. Our results show that the accuracy of TTT predictions depends sensitively on the smoothing scale used to compute the tidal field. At fixed smoothing lengths, smaller scales yield more accurate predictions for both the amplitude and orientation of the AM, particularly for low-mass protohalos. However, the optimal filter scale increases with protohalo mass, reflecting the coupling of larger protohalos to longer-wavelength tidal modes. Crucially, among the three protohalo size measures tested ($R_{\mathrm{eq}}$, $R_{\mathrm{lag}}$ and $10R_{\mathrm{vir}}$), the model achieves its best overall performance when the the filter scale is adjusted to the equivalent radius $R_{\mathrm{eq}}$ of each protohalo.
@@
In terms of direction, our results demonstrate a systematic and mass-dependent evolution of the AM orientation relative to the cosmic web. Early alignments with the spine of present-day filaments ($\hat{\mathbf{e}}_{3}$) degrade over time, particularly due to a gradual rotation of the AM vector around the intermediate filament axis ($\hat{\mathbf{e}}_{2}$), effectively transferring alignment toward the principal axis of collapse ($\hat{\mathbf{e}}_{1}$). This transition is most evident during the quasi-linear regime ($z\sim 10$ to $z\sim 1$), after which the highly non-linear regime tends to further degrade AM coherence. In this stage, the evolution can involve genuinely chaotic trajectories in some systems (e.g. due to the interaction of multiple halos), but in other cases it may follow more predictable, non-linear dynamics that are not captured by TTT. Despite this degradation, some coherence persists, especially for massive halos. Notably, the final alignments at $z=0$ appear more strongly shaped by the quasi-linear evolution than by the later transformations of the highly non-linear stage. These findings suggest that the breakdown of TTT is not abrupt but unfolds progressively, driven by the combined effect of internal dynamical processes, the differential coupling with the large-scale tidal field, and the evolving structure of the cosmic web.
@@
Altogether, our results show that both the depth of particles within halos and the time evolution of the protohalo shape critically influence the AM direction. The evolution from the initial conditions, as predicted by the standard TTT, to the stage we refer to as “highly non-linear” is not purely stochastic. Instead, it shows a strong correlation with the direction of the cosmic web, and in the quasi-linear regime this evolution appears to be statistically predictable when the time dependence of the inertia tensor is considered, as shown in Section 6.4. This predictable component suggests that part of the deviations from standard TTT can be understood in terms of systematic shape changes rather than random, late-time processes.
In this picture, early quasi-linear torques can even generate strong alignments even for initially weakly aligned systems (as we see, for instance, in Figure 10, between $\hat{\mathbf{J}}$ and $\hat{\mathbf{e}}_{3}$ for high-mass halos). Although the final stages of evolution can partially dilute these alignments, the $z=0$ signal still appears to reflect more strongly the imprint of the quasi-linear evolution than the influence of highly non-linear effects. The fact that a substantial fraction of the alignment evolution occurs in a regime where shape changes are measurable and predictable opens the possibility of improving analytical models without the need to explicitly incorporate all the details of highly non-linear dynamics. Properly modeling the inertia tensor evolution during this phase is thus essential to extend the predictive range of TTT beyond the initial linear regime.
In this context, the modified TTT formulation proposed by López & Merchán (2023) provides an interesting and complementary perspective. Their approach introduces anisotropic scale factors as a proxy for the influence of the large-scale tidal field, which in practice leads to systematic and predictable changes in halo shapes. While their model does not explicitly track the time evolution of the inertia tensor, it may nevertheless capture some of the same physical effects identified in our analysis, namely, that controlled and measurable shape variations can play a decisive role in setting the final ($z=0$) AM alignments. This connection suggests that further exploration of shape-driven extensions to TTT could provide a practical path toward more accurate and physically motivated predictions for halo AM correlations.
@@ Acknowledgements
This work has been partially supported by the Argentine Consejo Nacional de Investigaciones Científicas y Técnicas (CONICET). This work used computational resources from UNC Supercómputo (CCAD) – Universidad Nacional de Córdoba (https://supercomputo.unc.edu.ar), which are part of SNCAD, República Argentina. We also thank the referee Mark Neyrinck for his insightful comments, which have helped to significantly improve this paper.
@@ Appendix A Cosmic web identification with NEXUS+
The NEXUS+ suite of cosmic web identifiers represents an elaboration and extension of the original Multiscale Morphology Filter (Aragón-Calvo et al., 2007b, a, 2010) algorithm and was developed with the goal of obtaining a more physically motivated and robust method (for short summaries see also Libeskind et al. (2018); Kugel & van de Weygaert (2024)). NEXUS+ is the principal representative of the full NEXUS+ suite of cosmic web identifiers (see Cautun et al., 2013). These include the options for multiscale morphology identifiers on the basis of the linear density, the logarithmic density, the velocity divergence, the velocity shear and tidal force field. NEXUS+ has incorporated these options in a versatile code for the analysis of cosmic web structure and dynamics following the realization that they are significant physical influences in shaping the cosmic mass distribution into the complexity of the cosmic web.
@@ A.1 Hessian Geometry and Morphological Identity
The basic setup of MMF/NEXUS+ is that of defining a four-dimensional scale-space representation of the input field $f(\vec{x})$. In nearly all implementations this achieved by means of a Gaussian filtering of $f(\vec{x})$ over a set of scales $R_{0},R_{1},...,R_{N}]$,
@@
$H_{ij,R_{n}}(\vec{x})$ $\,=\,$ $R_{n}^{2}\;\frac{\partial^{2}f_{R_{n}}(\vec{x})}{\partial x_{i}\partial x_{j}}\,.\$(10)
$\hat{H}_{ij,R_{n}}(\vec{k})$ $\,=\,$ $-k_{i}k_{j}R_{n}^{2}\hat{f}(\vec{k})e^{-k^{2}R_{n}^{2}/2}\,.$
Note that the definition for the Hessian includes the normalization term $R_{n}^{2}$. The key element of the MMF/NEXUS+ formalism is the morphological information contained in the eigenvalues of the Hessian matrix, $h_{1}\leq h_{2}\leq h_{3}$. By applying a set of morphology filters on these scaled eigenvalues (see Aragón-Calvo et al., 2007a; Cautun et al., 2013) this is translated into a scale dependent environment signature $\mathcal{S}_{R_{n}}(\mathbf{x})$ that represents the geometry at the corresponding scale.
@@ A.2 Scale Space and Multiscale Structure
To analyse the multiscale nature of the cosmic web, the Scale-Space representation of the cosmic mass distributions produces a sequence of copies of the data having different resolutions (Florack et al., 1992; Lindeberg, 1994). At each location $\vec{x}$ in the probed volume, it involves an extra dimension, scale, that yields the eigenvalues of the Hessian filtered at the corresponding (Gaussian) scale and the scale-dependent environment signature $\mathcal{S}_{R_{n}}(\mathbf{x})$.
A feature searching algorithm is applied to the combined set of scaled copies in order to identify the scale at which, locally, we find the strongest morphological signature. It involves the combination of the complete set of scale-dependent environmental signatures to find the maximum signature for all scales:
@@ A.3 Signature & Versions
The final outcome of the MMF/NEXUS+ procedure is a field which at each location $\vec{x}$ specifies what the local morphological signature is, cluster node, filaments, wall or void. The resulting field $\delta^{\text{NEXUS}}_{j}(\mathbf{x})$ is zero when the volume is not identified as cosmic web element $j$ and is one when the volume element is identified as element $j$. Here $j$ is either filaments, nodes or walls. In this identification we also intrinsically include the identification for voids which is defined as the volume elements that are neither a filament, node or wall.
@@ A.4 NEXUS+
[…]
@@
The NEXUS+ algorithm performs the environment detection by applying the above steps first to knots, then to filaments and finally to walls. Each volume element is assigned a single environment characteristic by requiring that filament regions cannot be knots and that walls regions cannot be either knots or filaments. The remaining regions are classified as voids. An example of this classification, showing the density field together with the corresponding NEXUS+ signatures for nodes, filaments and walls, is presented in Figure 13.